OpenStudy (anonymous):

suppose you had d dollars in your bank account you soent 19$ but have at least $62 left. how much money did you have intitially. write and solve an inequality that represents this situation

OpenStudy (anonymous):

hmm .. isnt it simple addition and subtraction?

OpenStudy (anonymous):

x-19=62

OpenStudy (anonymous):

19+62=81 dollars the equation is x-19=62

OpenStudy (anonymous):

where x is the "d"